# Reverse Diffusion Sequential Monte Carlo Samplers

### v1

@@ 3.1 Reverse diffusion proposal
This sampling procedure builds on the prior work of Huang et al. [8], Grenioux et al. [9], which estimates the score function using MCMC. In contrast, we specifically adopt an IS- or SMC-based score estimate that also provides an unbiased estimate of the normalization constant. This property will be leveraged in the next section to construct intermediate target distributions for SMC.
@@ 3.2 Intermediate targets
Samples from the reverse diffusion proposal deviate from the true target due to two sources of error: time discretization of the reverse SDE, and bias in score estimation. To correct for these discrepancies, we design a series of intermediate target distributions. We first characterize the optimal but intractable targets, and then develop practical approximations using byproducts of score estimation. Finally, we extend the intermediate targets to include auxiliary variables. Our construction aligns with the ideal targets at intermediate steps and recovers the desired marginal $\pi(x_{0})$ at the final step $t=0$.
@@
The optimal intermediate target at step $t$ is the marginal distribution of the extended target $\pi(x_{0:T})$ defined in Eq.˜4:
$\gamma_{t}^{*}(x_{t:T})=\int\pi(x_{0:T})dx_{0:t-1}=p(x_{t})\prod_{i=t+1}^{T}p(x_{i}\mid x_{i-1}).$ (9)
This choice is optimal as it leads to exact samples from $\pi(x_{0})$ with locally optimal proposals [22].
However, the marginal $p(x_{t})=\int\pi(x_{0})\mathcal{N}(x_{t}\mid\alpha_{t}x_{0},\sigma_{t}^{2}\mathbb{I})dx_{0}$ in Eq.˜9 involves an intractable integral, except in the final step $t=0$, where $p(x_{0})=\pi(x_{0})$ is known up to a normalization constant.
Approximating the marginal.
At step $t>0$, we approximate the marginal distribution $p(x_{t})$ by $\hat{p}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$ using the same auxiliary variables $\color[rgb]{0.3125,0.46875,0.62890625}u_{t}$} used for the score estimation. The key insight is that $Zp(x_{t})$ is the normalization constant of the posterior $p(x_{0}\mid x_{t})\propto\tilde{\pi}(x_{0})\mathcal{N}(x_{t}\mid\alpha_{t}x_{0},\sigma_{t}^{2}\mathbb{I})$. Thus, by performing IS (or a similar AIS / SMC procedure) targeting $p(x_{0}\mid x_{t})$, we obtain importance weights $\{w^{(m)}\}_{m=1}^{M}$ that provide an unbiased estimate of $p(x_{t})$ up to the normalization constant $Z$. We define
$\hat{p}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$ $:=\frac{1}{M}\sum_{m=1}^{M}w^{(m)},\textrm{ where we have }\mathbb{E}_{q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\mid x_{t})}\left[\hat{p}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\right]=Zp(x_{t}).$ (10)
At the final step $t=0$, we set $\hat{p}(x_{0},{\color[rgb]{0.3125,0.46875,0.62890625}u_{0}}):=\tilde{\pi}(x_{0})=Z\pi(x_{0})$ as no approximation is needed.
See Algorithm˜2 for an example of utilizing the earlier IS score estimator to compute $\hat{p}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$.
@@
Finally, we incorporate the auxiliary variables by setting the targets to match their sampling distributions and define the intermediate target over $\{x_{t:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}}\}$ as
$\gamma_{t}(x_{t:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}})$ $:=\hat{p}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\,q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\mid x_{t})\prod_{i=t}^{T-1}p(x_{i+1}\mid x_{i})\,q({\color[rgb]{0.3125,0.46875,0.62890625}u_{i+1}}\mid x_{i+1})$ (11)
for $t=0,\cdots,T-1$, and $\gamma_{T}(x_{T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{T}}):=\hat{p}(x_{T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{T}})\,q({\color[rgb]{0.3125,0.46875,0.62890625}u_{T}}\mid x_{T})$.
@@
The structure in Eq.˜11 mirrors the optimal target in Eq.˜9, replacing the intractable $p(x_{t})$ with an unbiased estimate $\hat{p}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$ (up to a factor of $Z$) and accounting for auxiliary randomness. In fact, marginalizing out the auxiliary variables $\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}$} in $\gamma_{t}(x_{t:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}})$ recovers the optimal target
$\begin{split}&\int\gamma_{t}(x_{t:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}})d{\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}}=\mathbb{E}_{q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\mid x_{t})}\left[\hat{p}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\right]\,\prod_{i=t}^{T-1}p(x_{i+1}\mid x_{i})\\ &=Zp(x_{t})\,\prod_{i=t}^{T-1}p(x_{i+1}\mid x_{i})\propto\gamma_{t}^{*}(x_{t:T}),\end{split}$ (12)
where the last line follows from the unbiasedness property in Eq.˜10, for any step $t>0$. This is often referred to as an exact approximation [21, 22], as we still target the true $\gamma_{t}^{*}$ despite the approximation.
Crucially, the marginal of the final target $\gamma_{0}(x_{0})$ matches the desired target $\pi(x_{0})$,
@@
$\gamma_{0}(x_{0})$ $=\int\tilde{\pi}(x_{0})\prod_{i=1}^{T}p(x_{i}\mid x_{i-1})\,q({\color[rgb]{0.3125,0.46875,0.62890625}u_{i}}\mid x_{i})dx_{1:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{0:T}}\,\propto\pi(x_{0}),$ (13)
as $\hat{p}(x_{0},{\color[rgb]{0.3125,0.46875,0.62890625}u_{0}})=\tilde{\pi}(x_{0})$ is exact by construction.
The observations from Eqs.˜12 and 13 ensure that the SMC procedure we describe next provides effective guidance at intermediate steps while correctly targeting the final distribution.
@@ 3.3 RDSMC: Algorithm and theoretical guarantee
Input: Unnormalized target $\tilde{\pi}(x_{0})$, number of particles $N$, discretization steps $T$, score estimation budget $M$, diffusion schedule $\{\alpha_{t},\sigma_{t},f_{t},g_{t}\}$ and reference distribution $p(x_{T})$
@@
3   Sample $x_{T}^{(i)}\sim p(x_{T})$;
4   Compute score and marginal estimates $s_{T}^{(i)},\hat{p}_{T}^{(i)}\leftarrow\text{Algorithm 2}(\pi,\alpha_{T},\sigma_{T},x_{T}^{(i)},M)$;
5   Compute weight: $w_{T}^{(i)}\leftarrow\frac{\hat{p}_{T}^{(i)}}{p(x_{T}^{(i)})}$;
@@
9   Resample $\{x_{t+1}^{(i)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}}^{(i)},s_{t+1}^{(i)},\tilde{\pi}_{t+1}^{(i)}\}_{i=1}^{N}$ according to weights $\{w_{t+1}^{(i)}\}_{i=1}^{N}$ ;
@@
13        Compute score and marginal estimates $s_{t}^{(i)},\hat{p}_{t}^{(i)}\leftarrow\text{Algorithm 2}(\pi,\alpha_{t},\sigma_{t},x_{t}^{(i)},M)$;
@@
15        Compute exact marginal $\hat{p}_{0}^{(i)}\leftarrow\tilde{\pi}(x_{0}^{(i)})$;
@@
17    Compute weight: $w_{t}^{(i)}\leftarrow\frac{\hat{p}_{t}^{(i)}p(x_{t+1}^{(i)}\mid x_{t}^{(i)})}{\hat{p}_{t+1}^{(i)}q(x_{t}^{(i)}\mid x_{t+1}^{(i)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}}^{(i)})}$;
@@
With the proposals and intermediate targets in place, we now derive tractable weighting functions. Finally, we present the complete RDSMC algorithm and establish its theoretical guarantees.
@@
For $t=T-1,\cdots,0$, the intermediate weight is computed as:
$w_{t}$ $=\frac{\gamma_{t}(x_{t:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}})}{\gamma_{t+1}(x_{t+1:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1:T}})\,q(x_{t}\mid x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})\,q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\mid x_{t})},$ (14)
substituting intermediate target expressions from Eq.˜11 and canceling out common terms, we obtain
$w_{t}$ $=\frac{\hat{p}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})p(x_{t+1}|x_{t})}{\hat{p}(x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})q(x_{t}\mid x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})}.$ (15)
At the initial step $T$, the weight is $w_{T}=\frac{\gamma_{T}(x_{T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{T}})}{q(x_{T})q({\color[rgb]{0.3125,0.46875,0.62890625}u_{T}}\mid x_{T})}=\frac{\hat{p}(x_{T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{T}})}{p(x_{T})}$.
Notably, although the intermediate proposals and targets involve auxiliary sampling distributions, the weighting functions remain tractable and agnostic to them. This result allows the use of a fairly generic auxiliary sampler for score and marginal estimation, while still yielding computable weights.
@@
We summarize the RDSMC algorithm in Algorithm˜1. It initializes $N$ particles from the reference distribution. Subsequently, the particles are propagated through a diffusion-based proposal, followed by reweighting and resampling using intermediate targets. These steps are enabled by MC sampling from the posterior $p(x_{0}\mid x_{t})$. The final output is a weighted set of samples approximating the target $\pi(x)$, along with an estimate of the normalization constant $Z$. While Algorithm˜1 uses IS from Algorithm˜2 for MC sampling, RDSMC can incorporate more advanced AIS or SMC methods. See §˜B.1 for details.
RDSMC can be viewed as an adaptation of the nested SMC framework [23, 28]. Specifically, each particle at step $t$ involves running an inner-level MC estimator to compute the score function and the marginal density. This inner estimator constitutes an IS or a similar procedure that targets the posterior distribution $p(x_{0}\mid x_{t})$, which is then used to assign proper weights [22, Chapter 4.3] to proposed samples $x_{t}$. Consequently, the proposed RDSMC framework naturally inherits the unbiasedness and asymptotic exactness guarantees of such SMC algorithms [23, 22].
Theorem 1.
The RDSMC algorithm provides a consistent estimator of the target distribution $\pi(x_{0})$ as particle size $N\to\infty$ and an unbiased estimator of the normalization constant $Z$ for any $N\geq 1$.
We provide the formal statement and proof in Appendix˜C, under standard regularity assumptions.

### v2

@@ 3.1 Reverse diffusion proposal
This sampling procedure builds on prior work of huang2023reverse, grenioux2024stochastic, which estimate the score function using MCMC, and by he2024zeroth which use rejection sampling. We formalize these ideas by defining an extended proposal distribution that incorporates randomness in MC score estimates. In particular, we adopt an IS or AIS-based MC approach that also provides an unbiased estimate of the normalization constant, a property that, in principle, can be achieved by the rejection sampling approach of he2024zeroth as well. This property will be leveraged in the next part to construct intermediate targets for RDSMC.
@@ 3.2 Intermediate targets
Samples from the reverse diffusion proposal deviate from the desired target $\pi$ due to two sources of error: time discretization of the reverse SDE, and bias in score estimation. To correct these errors, we design a series of intermediate target distributions. We first characterize the optimal but intractable targets, and then develop practical approximations using byproducts of score estimation. Finally, we extend the intermediate targets to include auxiliary variables. Our construction aligns with the optimal targets at intermediate steps and recovers the desired marginal $\pi(x_{0})$ at the final step $t=0$.
@@
The optimal intermediate target at step $t$ is the marginal distribution of the extended target $\pi(x_{0:T})$ in Eq.˜4,
$\pi(x_{t:T})=\int\pi(x_{0:T})\mathrm{d}x_{0:t-1}=\pi(x_{t})\prod_{i=t+1}^{T}\pi(x_{i}\mid x_{i-1}),$ (9)
as it leads to exact samples from $\pi(x_{0:T})$ when combined with locally optimal proposals [naesseth2019elements].
However, the marginal $\pi(x_{t})=\int\pi(x_{0})\mathcal{N}(x_{t}\mid\alpha_{t}x_{0},\sigma_{t}^{2}\mathbb{I})\mathrm{d}x_{0}$ in Eq.˜9 involves an intractable integral, except in the final step $t=0$, where $\pi(x_{0})$ is known up to a normalization constant.
Marginal estimation.
For $t>0$, we approximate the marginal $\pi(x_{t})$ by $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$ using byproducts of MC score estimates, where we recall $\color[rgb]{0.3125,0.46875,0.62890625}u_{t}$} is the estimation randomness. The key insight is that $Z\pi(x_{t})$ is the normalization constant of the posterior $\pi(x_{0}\mid x_{t})\propto\tilde{\pi}(x_{0})\mathcal{N}(x_{t}\mid\alpha_{t}x_{0},\sigma_{t}^{2}\mathbb{I})$, that is,
$\int\tilde{\pi}(x_{0})\mathcal{N}(x_{t}\mid\alpha_{t}x_{0},\sigma_{t}^{2}\mathbb{I})\mathrm{d}x_{0}$ $=Z\int\pi(x_{0})\mathcal{N}(x_{t}\mid\alpha_{t}x_{0},\sigma_{t}^{2}\mathbb{I})\mathrm{d}x_{0}=Z\pi(x_{t}).$
Hence, we can re-use the posterior inference procedure in MC score estimation to obtain an unbiased estimate of the desired marginal $\pi(x_{t})$ (up to a factor of $Z$).
For example, IS or AIS-based score estimation (Algorithms˜2 and 3) generates importance weights $\{w^{(m)}\}_{m=1}^{M}$ targeting $\pi(x_{0}\mid x_{t})$ for a fixed $x_{t}$; we then obtain an unbiased marginal estimate
$\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$ $\coloneqq\frac{1}{M}\sum_{m=1}^{M}w^{(m)},\textrm{ with }\mathbb{E}_{q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\mid x_{t})}\left[\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\right]=Z\pi(x_{t}).$ (10)
At the final step $t=0$, we set $\hat{\pi}(x_{0},{\color[rgb]{0.3125,0.46875,0.62890625}u_{0}})\coloneqq\tilde{\pi}(x_{0})=Z\pi(x_{0})$ as no approximation is needed.
@@
Finally, we incorporate the auxiliary variables $\color[rgb]{0.3125,0.46875,0.62890625}u_{t}$} by setting their targets to match their sampling distributions and define the extended intermediate target as
$\gamma_{t}(x_{t:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}})$ $\coloneqq\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\,q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\mid x_{t})\prod_{i=t}^{T-1}\pi(x_{i+1}\mid x_{i})\,q({\color[rgb]{0.3125,0.46875,0.62890625}u_{i+1}}\mid x_{i+1})$ (11)
for $t=0,\cdots,T-1$, and $\gamma_{T}(x_{T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{T}})\coloneqq\hat{\pi}(x_{T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{T}})\,q({\color[rgb]{0.3125,0.46875,0.62890625}u_{T}}\mid x_{T})$.
@@
The structure in Eq.˜11 mirrors that of the optimal targets in Eq.˜9, replacing the intractable marginal $\pi(x_{t})$ with an unbiased estimate $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$ (up to a factor of $Z$) and accounting for auxiliary randomness $\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}$}. Moreover, we make the following two observations.
Observation 1.
The final marginal target at $t=0$ matches the desired $\pi(x_{0})$ with no approximations,
@@
$\begin{split}\gamma_{0}(x_{0})&=\int\gamma_{0}(x_{0:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{0:T}})\mathrm{d}x_{1:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{0:T}}\\ &=\int\hat{\pi}(x_{0},{\color[rgb]{0.3125,0.46875,0.62890625}u_{0}})q({\color[rgb]{0.3125,0.46875,0.62890625}u_{0}}\mid x_{0})\prod_{i=1}^{T}\pi(x_{i}\mid x_{i-1})\,q({\color[rgb]{0.3125,0.46875,0.62890625}u_{i}}\mid x_{i})\mathrm{d}x_{1:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{0:T}}\\ &=\tilde{\pi}(x_{0})\propto\pi(x_{0}),\end{split}$ (12)
as $\hat{\pi}(x_{0},{\color[rgb]{0.3125,0.46875,0.62890625}u_{0}})=\tilde{\pi}(x_{0})$ by construction. This implies our final SMC target is correctly specified.
Observation 2.
Marginalizing out $\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}$} in $\gamma_{t}(x_{t:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}})$ recovers the optimal target $\pi(x_{t:T}),\forall t$,
$\begin{split}\int\gamma_{t}(x_{t:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}})\mathrm{d}{\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}}&=\int\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\,q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\mid x_{t})\prod_{i=t}^{T-1}\pi(x_{i+1}\mid x_{i})\,q({\color[rgb]{0.3125,0.46875,0.62890625}u_{i+1}}\mid x_{i+1})\mathrm{d}{\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}}\\ &=\mathbb{E}_{q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\mid x_{t})}\left[\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\right]\,\prod_{i=t}^{T-1}\pi(x_{i+1}\mid x_{i})\\ &=Z\pi(x_{t})\,\prod_{i=t}^{T-1}\pi(x_{i+1}\mid x_{i})\propto\pi(x_{t:T}),\end{split}$ (13)
where the last equality follows from the unbiasedness of marginal estimates, as in Eq.˜10. This result shows that our intermediate targets $\gamma_{t}(x_{t:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}})$ are aligned with the optimal $\pi(x_{t:T})$, despite the approximation in $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$, a property known as the exact approximation [andrieu2010particle, naesseth2019elements].
While the generic SMC framework remains valid under alternative choices of intermediate targets, provided that $\gamma_{0}(x_{0})\propto\pi(x_{0})$, our construction offers a balance between theoretical correctness and practical efficiency. The marginal estimates $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$ in our intermediate targets from Eq.˜11 act as twisting or look-ahead functions [naesseth2019elements, Chapter 3] that incorporate future information $\pi(x_{0})$ into intermediate steps. Together, Observations˜1 and 2 ensure that our intermediate targets provide effective guidance throughout the sampling process while recovering the desired target in the end.
@@ 3.3 RDSMC: Algorithm and theoretical guarantee
We now derive weighting functions, the final component of RDSMC, which guides the resampling of particles. We then present the complete algorithm and establish its theoretical guarantees.
Input: Unnormalized target $\tilde{\pi}(x_{0})$, number of particles $N$, discretization steps $T$ (with step size $\delta=1/T$), diffusion schedule $\{\alpha_{t},\sigma_{t},f_{t},g_{t}\}$, base distribution $q(x_{T})$, and additional inputs for score and marginal estimation $\eta$
@@
3   Sample $x_{T}^{(i)}\sim q(x_{T})$;
[…]
5   Compute weight: $w_{T}^{(i)}\leftarrow\frac{\hat{\pi}_{T}^{(i)}}{q(x_{T}^{(i)})}$;
@@
9   Resample $\{x_{t+1}^{(i)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}}^{(i)},s_{t+1}^{(i)},\hat{\pi}_{t+1}^{(i)}\}_{i=1}^{N}$ according to weights $\{w_{t+1}^{(i)}\}_{i=1}^{N}$ ;
@@
13        Compute score and marginal estimates $s_{t}^{(i)},\hat{\pi}_{t}^{(i)}\leftarrow$ Algorithm 3 $(\pi,\alpha_{t},\sigma_{t},x_{t}^{(i)},\eta)$;
@@
15        Compute exact marginal $\hat{\pi}_{0}^{(i)}\leftarrow\tilde{\pi}(x_{0}^{(i)})$;
@@
17    Compute weight: $w_{t}^{(i)}\leftarrow\frac{\hat{\pi}_{t}^{(i)}\,\pi(x_{t+1}^{(i)}\mid x_{t}^{(i)})}{\hat{\pi}_{t+1}^{(i)}\,q(x_{t}^{(i)}\mid x_{t+1}^{(i)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}}^{(i)})}$;
@@
Following the standard SMC framework, we define the weighting functions on the extended space $\{x_{t:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}}\}$ for $t=T-1,\cdots,0$ as ratios of intermediate targets and proposals
$w_{t}$ $\coloneqq\frac{\gamma_{t}(x_{t:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}})}{\gamma_{t+1}(x_{t+1:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1:T}})\,q(x_{t}\mid x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})\,q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\mid x_{t})}.$ (14)
Substituting the expressions for $\gamma_{t}$ from Eq.˜11 and canceling out common terms yields
$w_{t}$ $=\frac{\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\pi(x_{t+1}|x_{t})}{\hat{\pi}(x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})q(x_{t}\mid x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})}.$ (15)
At the initial step $T$, the weight is $w_{T}\coloneqq\frac{\gamma_{T}(x_{T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{T}})}{q(x_{T})q({\color[rgb]{0.3125,0.46875,0.62890625}u_{T}}\mid x_{T})}=\frac{\hat{\pi}(x_{T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{T}})}{q(x_{T})}$ recalling that $\gamma_{T}(x_{T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{T}})=\hat{\pi}(x_{T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{T}})q({\color[rgb]{0.3125,0.46875,0.62890625}u_{T}}\mid x_{T})$ by construction.
Notably, although the intermediate targets and proposals involve auxiliary sampling distributions $q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\mid x_{t})$ for score estimation, the weighting functions $w_{t}$ in Eq.˜15 and $w_{T}$ do not depend on their evaluation. Hence, a generic MC sampler can be used while still retaining computable weights, as long as it produces tractable estimates of the score and marginal densities.
@@
We summarize the RDSMC algorithm in Algorithm˜1. It initializes $N$ particles from a tractable base distribution. Subsequently, the particles are propagated through a reverse diffusion-based proposal, followed by weighting and resampling using intermediate targets. These steps are enabled by running an inner-level MC estimation of the score function and the marginal density. In this work, we use AIS (Algorithm˜3) while other methods can be incorporated as well (see §˜B.1). The final output of RDSMC is a weighted set of samples approximating the target $\pi(x_{0})$, along with an estimate of the normalization constant $Z$.
RDSMC can be viewed as an adaptation of the nested SMC framework [naesseth2015nested, naesseth2019high]. The inner-level MC estimation involves a sampling procedure (e.g. AIS) targeting the intractable posterior $\pi(x_{0}\mid x_{t})$, which is then used to assign proper weights [naesseth2019elements, Chapter 4.3] to proposed samples $x_{t}$. Consequently, RDSMC inherits the unbiasedness and asymptotic exactness guarantees of nested SMC [naesseth2015nested, naesseth2019elements].
Theorem 1 (Informal).
Under regularity conditions, the RDSMC algorithm provides a consistent estimator of the target distribution $\pi(x_{0})$ as particle size $N\to\infty$ and an unbiased estimator of the normalization constant $Z$ for any $N\geq 1$.
We provide the formal statement and proof in Appendix˜C.
Theorem˜1 shows that RDSMC effectively wraps a diffusion MC-style proposal within an SMC framework, mitigating its bias as the number of particles increases. Moreover, it provides an unbiased estimate of the normalization constant, a property absent in existing diffusion MC methods [huang2023reverse, grenioux2024stochastic, he2024zeroth].
The bias correction mechanism of RDSMC arises from two key aspects: (1) the final target of RDSMC is explicitly constructed to match the desired target $\pi(x_{0})$, regardless of the discretization scheme or marginal density approximation (Observation˜1); and (2) errors in score estimation affect only the proposal steps, which are accounted for in the weighting functions. Therefore, the weighting and resampling steps in the outer SMC loop asymptotically correct for proposal bias, ensuring convergence to the final target $\pi(x_{0})$ in the limit of many particles.